STRUCTURE AND DYNAMICS OF THE CYCLOPROPENE RADICAL CATION STUDIED BY HIGH-RESOLUTION PHOTOELECTRON SPECTROSCOPY

\underline{KONSTANTINA VASILATOU}, JULIE M. MICHAUD, DENITSA BAYKUSHEVA, GUIDO GRASSI and FREDERIC MERKT, Laboratorium für Physikalische Chemie, ETH Zürich, CH-8093 Zürich, Switzerland.

High-resolution photoelectron spectra of cyclopropene (c-C_{3}H_{4}) and several of its deuterated isotopomers have been recorded. From the analysis of the almost fully resolved rotational structure of the origin of the X̃^{+} ^{2}B_{2} ← X̃^{1}A_{1} ionizing transition of c-C_{3}H_{4}, c-C_{3}H_{3}D, c-C_{3}H_{2}D_{2}, c-C_{3}HD_{3} and c-C_{3}D_{4} and the observation of several low-lying vibrational levels of the corresponding cations, we conclude that the cyclopropene cation has an equilibrium structure deviating from C_{2v} symmetry. We believe that the cation is subject to a large-amplitude tunnelling motion along a mode of a_{2} symmetry involving a twist of the CH_{2} group and an asymmetric out-of-plane bending of the CH units. The reasons for this belief and also currently unresolved aspects of our analysis will be discussed.